I am not certain if I need to use the a^ +b^ =c and then use a radical expression to reduce.
One side of a rectangular stage is
2 meters longer than the other. If the diagonal is 10 meters,
then what are the lengths of the sides?
Thank you,
Tracy

Answer by ankor@dixie-net.com(12701)

One side of a rectangular stage is 2 meters longer than the other.
If the diagonal is 10 meters, then what are the lengths of the sides?
:
Actually it's:
a^2 + b^2 = c^2
:
In this problem call one side x:
Let x = a
then
(x+2) = b
and
c = 10
:
Substituting in our formula:
x^2 + (x+2)^2 = 10^2
:
FOIL (x+2)(x+2)
x^2 + (x^2 + 4x + 4) = 100
:
Combine like terms, arrange as a quadratic equation
x^2 + x^2 + 4x + 4 - 100 = 0
2x^2 + 4x - 96 = 0
:
Factor
(2x - 12)(x + 8) = 0
;
We want the positive solution here
2x = 12
x = 6 meters is the 1st side
then
6 + 2 = 8 meters is the 2nd side
:
:
We can check that: 6^2 + 8^2 = 10^2
:
Was this understandable? Any questions?